# Non-Gaussian enhancement of precision in quantum batteries

### v1

@@ Abstract
In Fig.4 we show the dependence on temperature of the highest ratio that can be attained for a different number of ladder operations, after optimal tuning of the Gaussian block’s $U_{G}$ parameters. In addition, the behavior of $\Gamma$ is also shown for the superposition of a coherent and a Fock (with $n=5$) state in contact with a thermal bath at temperature $T$, whose photon-number statistical moments are derived in Weber et al. (2022), and a noisy kitten state, modeled as a photon-subtracted squeezed thermal state, as suggested in Ourjoumtsev et al. (2006); Wakui et al. (2007). The analytics for the calculation of $\Gamma$ for the kitten states can be found in the SM.
Different kinds of non-Gaussian states can violate the Gaussian bound (black dashed line). Photon-added states become Fock states at zero temperature, and the SNR explodes as their fluctuations vanish. Noisy cat states showcase a similar behaviour, although their advantage with respect to the Gaussian bound is more sensitive to fluctuations than it is for photon-added states. On the contrary, coherent-Fock thermal states appear to be more noise-robust and display more stable signal-to-noise ratios, violating the Gaussian bound in the low thermal fluctuations regime. We also observe that photon-added states are more resilient to noise if we increase the number of photon additions. The intersections of the solid lines in Fig.4 with the Gaussian bound indicate the temperatures at which one more photon addition is needed to overcome the best Gaussian ratio. Whether these photon-added states can be considered as potential candidates for the charging process or not depends on the value of the mean photon number constraint $\epsilon$ imposed by each specific use case. A detailed discussion on the conditions for feasibility is provided in the SM.
Figure 4: Optimal $\Gamma$ attainable for zero (Gaussian) to three photon-added states, thermal-Fock and noisy cat states, as a function of temperature. The constraint has been set at $\epsilon=5$ (dimensionless).
Conclusion.– In this work, we describe the performance of bosonic quantum batteries under realistic physical constraints. We introduce a new figure of merit, the SNR of the energy conversion, that is practically motivated by the precision charging of nano-scale energy harvesters. This figure of merit naturally introduces a hierarchy of the type of states used as batteries. We find optimal values of the experimental parameters required to implement the batteries that minimize the fluctuations of the energy conversion. We find connections between the SNR and other relevant quantities in quantum optics and quantum information and use them to show that it can be an indicator of non-classicality and non-Gaussianity. Furthermore, we provide evidence of the potential gain in SNR through the application of entangling operations in the non-Gaussian realm. We provide analytical calculations to support our findings that can be easily adapted to other physically relevant scenarios. A possible follow-up of this work would be to introduce a more realistic noise model and use quantum control theory to minimize the fluctuations of the batteries dynamically. In addition, our model paves the way to study how non-Gaussianity affects fluctuations in the context of thermodynamic uncertainty relations. The experimental implementation of this protocol would constitute the first demonstration of quantum advantage for a thermodynamic practical scenario, with relevant applications in the miniaturization of nanotechnologies.
@@
Acknowledgements: The authors thank Antonis Delakouras, Manuel Gessner and Francesco Flora for fruitful discussions and inputs, and notably Antonio Acín for his relevant feedback and guidance.
@@
Funded by the European Union (EQC, 101149233), the Government of Spain (Severo Ochoa CEX2019-000910-S), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program). Views and opinions expressed are however, those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Appendix A Covariance matrix formalism for Gaussian states
@@
a squeezing transformation of parameter $z$. This representation allows for the interpretation of every Gaussian state as the result of applying a single mode squeezing transformation followed by a linear optical passive setup to an initial thermal (and possibly displaced) state of $N$ modes. Such initial state would have covariance matrix $\sigma_{0}=\bigoplus_{j=1}^{N}\nu_{j}\mathds{1}_{2}$, where the $\nu_{i}$ are referred to as symplectic eigenvalues, and can be understood in this context as fluctuation parameters indicating the temperature $T_{i}$ that characterizes the initial thermal state of each of the modes through the relation:
$\nu_{i}=\coth\bigg(\frac{\omega_{i}}{k_{B}\ T}\bigg{missing})$ (21)
The matrices $O_{1},O_{2}$ represent the energy-preserving passive optics transformations, which can be decomposed into beam splitters (semi-reflective mirrors parametrized by their transmissivity $\cos^{2}\theta$ that mix up pairs of modes) and phase shifters (dielectric plates that rotate the optical phase of an electromagnetic wave by some angle $\varphi$). A scheme of alternating applications of these two operations to create the most general linear interferometer is detailed in Tan and Rohde (2019). Matrix $Z$ represents local or single-mode squeezing, which, in contrast, is an active operation, requiring energy from an external laser field that pumps a nonlinear crystal through which the optical modes travel. The result of this transformation is a contraction of the variance of one of the canonical variables and the expansion of its conjugate one. Squeezing and beam splitting allow the generation of quantum correlations among different modes of the quantum state.
@@
$\sigma=\nu\begin{pmatrix}z\cos^{2}\varphi+\frac{1}{z}\sin^{2}\varphi&\frac{1-z^{2}}{2z}\sin(2\varphi)\\ \\ \frac{1-z^{2}}{2z}\sin(2\varphi)&zsin^{2}\varphi+\frac{1}{z}\cos^{2}\varphi\end{pmatrix}$ (22)
The mean and variance of the photon number operator can be obtained through (see Friis and Huber (2018))
@@
$\begin{split}(\Delta N(\rho))^{2}&=\frac{1}{8}\nu^{2}\bigg{(}z^{2}+\frac{1}{z^{2}}\bigg{)}-\frac{1}{4}\\ &+\nu\bigg{[}\alpha_{1}^{2}(zcos^{2}\varphi+\frac{1}{z}\sin^{2}\varphi)\\ &+\alpha_{2}^{2}(z\sin^{2}\varphi+\frac{1}{z}\cos^{2}\varphi)\\ &+2\alpha_{1}\alpha_{2}\bigg{(}\frac{1-z^{2}}{z}sin\varphi\ \cos\varphi\bigg{)}\bigg{]}\end{split}$ (26)
@@
and through the quotient between Eqs.25 and 28 one can compute the ratio $\Gamma$ of any Gaussian state as a function of the squeezing and displacement transformations.
Appendix B Optimal signal-to-noise ratio $\Gamma$ for Gaussian states
@@
and finally the optimal Gaussian ratios $\Gamma$, which are shown in Fig.2, are simply calculated by substituting these optimal solutions for $z$ and $\vec{\alpha}$ into Eqs.25 and 28.
To get a hint on how much of each energetic resource (squeezing and displacement) contributes to the optimal Gaussian state, we show the numerical values of the solutions as a function of $T,\epsilon$ in Fig.5. Whereas the optimal displacement seems to simply increase as more energy is allowed into the system, regardless of temperature, the optimal squeezing displays a less uniform behaviour. Interestingly, for most of the $(T,\epsilon)$ values, highly squeezed states (i.e. $z_{opt}$ small) are advantageous, and so for the majority of the parameter space we may approximate
@@
Figure 5: Solution to 29 for the optimal squeezing and displacement through the Lagrange multipliers method.
@@
Appendix C Expectation values on non-Gaussian states

### v2

@@ Abstract
In Fig. 4 we show, as a function of temperature, the highest ratio attainable for different numbers of ladder operations $m$, after optimal tuning of $U_{G}$. For comparison, we also plot $\Gamma$ for (i) superpositions of coherent and Fock states (with $n=5$) in contact with a thermal bath at temperature $T$ (moments from [WEBER2022168986]), and (ii) noisy kitten states modeled as photon-subtracted squeezed thermal states [kitten_states, Wakui2007]. Closed-form expressions for $\Gamma$ in the kitten case are provided in Appendix D.
Figure 4: Best-achieved SNR $\Gamma$ versus temperature for different non-Gaussian families (optimized over $U_{G}$), compared against the Gaussian bound (black dashed line). The energetic constraint has been set at $\epsilon=5$ (dimensionless).
Different non-Gaussian states violate the Gaussian bound (black dashed line). Photon-added states recover Fock states at $T=0$, with diverging SNR as fluctuations vanish; at finite temperature, their advantage persists up to a threshold set by $(\nu,\epsilon)$ and improves with increasing $m$. Noisy cat/kitten states display a similar low-temperature advantage but are more sensitive to thermal noise than photon-added states. Conversely, coherent–Fock thermal mixtures are comparatively noise-robust, yielding stable SNRs that still surpass the Gaussian bound in the low-$\nu$ regime. The intersections of the solid curves with the Gaussian bound indicate the temperatures at which one additional photon addition is required to maintain a non-Gaussian advantage. Whether these states are feasible candidates for the charging process depends on the mean-photon constraint $\epsilon$ relevant to the task; practical operating windows and constraints are discussed in Appendix E. We further show that, unlike the Gaussian case, passive interferometers can redistribute moments and strictly raise $\Gamma$ for non-Gaussian resources, enabling an interferometric optimization (see Appendix Figs. 78).
Within the fixed-energy setting $\delta N\leq\epsilon$, our results identify concrete, experimentally accessible non-Gaussian families that deliver a testable precision advantage for energy harvesting: observing $\Gamma>\Gamma_{G}(\nu,\epsilon)$ is both a performance gain and, by Theorem B, a witness of non-Gaussianity.
III.4 D. Multimode states
When the battery system comprises multiple modes, the possibility to perform entangling operations between them is unlocked. This naturally raises the question of whether such operations can offer an advantage in the relative precision of the stored energy [Gyhm_2024]. In the most general case, each of the modes may have different oscillating frequencies, and a comment in this regard is made in [Friis_2018] on the fact that Gaussian passive optics are sufficient to mix the modes in such a way that photons are stored in the lower frequency modes, implying a gain in energy precision.
However, in our scenario the response of the harvester is uniquely dependent on the photon number statistics (thus equivalent to the case where the battery modes all oscillate at the same frequency $\omega$). In that case, the following lemma applies:
Lemma 1.
For multimode Gaussian states, the signal-to-noise ratio $\Gamma$ remains constant under Gaussian global (entangling) passive transformations.
The proof is provided in Appendix G. On the contrary, non-Gaussian states feature signal-to-noise ratios that are tunable by passive, entangling operations, allowing for some extra advantage with respect to the Gaussian case when the correct beam-splitting angle is applied. In this sense, our figure-of-merit also constitutes a signature of non-Gaussianity in the multimode scenario, because the dependency of $\Gamma$ on mode mixing passive interferometry is an exclusive property of non-Gaussian states. In Appendix G we also discuss the scaling of the optimal $\Gamma$ with the number of modes, for Gaussian and non-Gaussian states of increasing number of photon additions, concluding that the non-Gaussian advantage is generalizable to the multi-mode scenario, and thus supporting the prospect of leveraging non-Gaussian resources for scalable and high-precision quantum batteries.
IV IV. Implementability
Our proposal is compatible with standard photonic hardware. The energy harvester can be a linear photodiode (Si or InGaAs PIN), whose average photocurrent is proportional to the absorbed photon flux. Writing $\eta$ for the overall quantum efficiency, $f$ for the pulse repetition rate, and $\langle N\rangle$ for the mean photons per pulse at the harvester, one has the usual relation
$\langle I\rangle\;=\;q\,\eta\,f\,\langle N\rangle\;=\;R\,P_{\mathrm{opt}},$ (18)
with $q$ the electron charge, $P_{\mathrm{opt}}$ the mean optical power, and $R=\eta q/(h\omega)$ the responsivity [RPPhotonics_Responsivity]. In what follows we assume the standard linear photodetection regime, in which the photodiode response is well described by normally ordered photocurrent moments for a single effective optical mode within the detection bandwidth [MandelWolf_OpticalCoherence, Loudon_QuantumTheoryLight]. Under this widely used model, the photocurrent statistics faithfully reproduce the photon-number statistics of the incident field up to an overall quantum efficiency $\eta$ and an additive electronic-noise contribution [MandelWolf_OpticalCoherence, Loudon_QuantumTheoryLight]. All witness statements based on electrical data should be understood as conditional on these linearity and calibration assumptions, which are routinely validated in precision photodetection experiments [LIGO_Squeezing_NatPhoton_2013, Berchera_2019, Ann2019].
Shot-noise-limited readout with commercial transimpedance/balanced receivers is routine when devices are chosen for low dark current and the front-end follows standard good practice [Analog_TIA_WideRange, Hamamatsu_KSPD9001E, Hamamatsu_PhotosensorModules_2025]. In this regime, with typical operation at $\langle N\rangle\sim 10^{3}$$10^{4}$ photons per detection bin at MHz repetition rates, pulse-to-pulse fluctuations of the harvested charge are directly resolved on photodiodes. In this regime, pulsed operation enables a per-pulse electrical access to non-classical photon-number statistics—for example, sub-shot-noise intensity fluctuations or $g^{(2)}(0)<1$—provided the linear-response assumptions above hold. Extensive prior art demonstrates that such non-classical signatures can be resolved using only photodiode readout in carefully calibrated setups [LIGO_Squeezing_NatPhoton_2013, Analog_TIA_WideRange, Ann2019, Berchera_2019]. By contrast, preparing genuinely non-Gaussian optical states at such large mean photon numbers is experimentally demanding: de-Gaussification via photon addition/subtraction is heralded and loss-sensitive, and non-Gaussian features tend to be diluted at high flux [walschaers2021non, kitten_states, Wakui2007]. Accordingly, we do not rely on bright non-Gaussian sources for our benchmark.
When the optical state is too dim for single-pulse resolution—e.g., $\langle N\rangle\lesssim 10$ photons per pulse—the same photodiode operates in an accumulate-and-reset mode with a charge-integrating front-end: the harvester integrates the photocurrent from $M$ pulses (typically $M\!\sim\!10^{4}$$10^{6}$, from ms to s depending on $f$) and the output voltage is sampled before reset. This replaces single-shot resolution with short-window statistics and is sufficient to observe the reduction of accumulated fluctuations. In this operation mode one can estimate $\Gamma_{\mathrm{det}}$ from the electrical data and compare it to the detection-propagated classical and Gaussian bounds (Sec. V.V and Appendix H H), thereby testing the implications of Theorems AB at the electrical level without modifying the optical source [TI_IVC102_Integrator, Hamamatsu_KSPD9001E, Hamamatsu_PhotosensorModules_2025]. The resulting certification of non-classicality or non-Gaussianity is therefore conditional on the same linear loss/noise model used to propagate the bounds. This division of labor—bright regime for non-classicality; accumulator regime for non-Gaussianity—matches the current state of the art on non-Gaussian state preparation, where demonstrated “kitten/cat” and photon-subtracted states typically have modest mean photon numbers and are fragile to loss [walschaers2021non, kitten_states, Wakui2007].
Our theoretical analysis is frequency-agnostic (we work with a number operator $N$ that neglects spectral structure). In practice, the field has a finite spectrum and the harvester/electronics have frequency responses. Implementations therefore account for (i) the photodiode spectral responsivity $R(\omega)$ and any optical filtering, which define an effective detected mode (or set of modes), and (ii) the electronic bandwidth of the readout relative to the pulse duration and repetition rate. These standard calibrations simply enter $\eta$ and $R$ and do not change the logic of the tests (non-classicality via $g^{(2)}(0)<1$; non-Gaussianity via $\Gamma>\Gamma_{G}$).
Looking beyond bulk photodiodes, nanoscale transducers offer a promising outlook for ultra-low-power loads. Colloidal quantum-dot photodiodes provide high responsivity and CMOS-compatible integration across the visible–SWIR bands [KonstantatosSargent_NatNano_2010], while resonator-coupled double quantum dots have demonstrated conversion of individual microwave photons into directed electrical current [QD_DoubleDot_Harvester_PRB_2024]. Although these platforms operate in spectrally distinct bands from our optical parameter choices, they reinforce that photon-to-charge transduction at ultra-low powers is feasible with state-of-the-art technology. They also suggest that, once suitable non-Gaussian sources are interfaced to such devices, our precision-charging bounds could be tested under realistically lossy conditions.
V V. Detectable precision
Theorems AB are state-level statements: $\Gamma=\delta N/\Delta N$ is a property of the photonic state delivered to the harvester and does not depend on how it is probed. In an experiment, however, we access a detection-level SNR after a standard linear loss channel of overall efficiency $\eta$ and additive, state-independent classical readout noise. As shown in Appendix H.1, such a lossy/noisy readout cannot increase SNR: the measured SNR $\Gamma_{\rm det}$ is always smaller or equal to the state SNR $\Gamma$ (equality only in the ideal, noiseless unit-efficiency limit). To retain the witness character when testing Theorems AB electrically, we therefore compare $\Gamma_{\rm det}$ to the detection-propagated classical and Gaussian bounds, obtained by pushing the corresponding state-level bounds through the same loss/noise map (see Appendix H.1).
Operationally, when we say that exceeding a detection-propagated Gaussian bound “certifies non-Gaussianity at the electrical readout”, we mean the following: if (i) the optical state delivered to the harvester is the only source of non-classical fluctuations, and (ii) the harvester plus electronics are well described by a linear loss channel of efficiency $\eta$ followed by additive, state-independent classical noise, then any observed $\Gamma_{\rm det}$ larger than the detection-propagated Gaussian bound cannot be reproduced by any Gaussian input state. Under these explicit assumptions, the electrical measurement functions as an indirect yet fully task-based witness of non-Gaussianity (Theorem B) and non-classicality (Theorem A).
Because the harvester ultimately dissipates power in a resistive element at (effective) temperature $T_{e}$, the finite-time thermodynamic uncertainty relation (TUR) imposes a detection-level ceiling on any integrated-current SNR over a window $\tau$: $\Gamma_{\tau}\leq\sqrt{\Sigma_{\tau}/2}$, where $\Sigma_{\tau}$ is the total entropy production in the load during $\tau$ [BaratoSeifert2015_TUR, Gingrich2016_DissipationBounds, DechantSasa2018_EntropicBound, KoyukSeifert2019_TimeDependentTUR, HorowitzGingrich2020_NatPhys_TURReview]. We use this bound as an absolute, dissipation-limited benchmark for the electrical readout (see Appendix H.2); it is independent of how the optical state is prepared.
As an illustration, for T=1,T=1,K, energy budget $\epsilon=5$, overall efficiency $\eta=0.60$, and electronic noise $\sigma_{B}=0.30$, a single photon–added state optimized under these constraints attains $\Gamma_{\mathrm{det}}\approx 2.21$, while the detection-propagated Gaussian bound is $\Gamma_{G,\mathrm{det}}\approx 1.84$ (Appendix H.1).
VI VI. Discussion and outlook
We analyze the precision charging for bosonic quantum batteries under a realistic finite-energy constraint and use the signal-to-noise ratio (SNR) of delivered excitations as an operational metric tied to what a harvester measures. On the theory side, we establish two bounds that transform this metric into witnesses: (i) a classical bound (Theorem A), certifying non-classicality, and (ii) a Gaussian bound (Theorem B) at fixed temperature and energy input, whose violation certifies non-Gaussianity. We then identify explicit, experimentally motivated non-Gaussian families that exceed the Gaussian bound in relevant parameter regimes and quantify how this advantage depends on temperature and energy budget. Conceptually, our results complement prior analyses of fluctuation-minimizing states in charging [Friis_2018] and studies using SNR in specific models [rinaldi2024reliablequantumadvantagequantum]: we provide a finite-energy Gaussian bound directly in the conversion metric measured at the load and show how concrete non-Gaussian resources surpass it.
On the implementation side, we describe a minimal, photodiode-based route that leverages standard hardware: a bright-pulse mode to access non-classical photon-number statistics per pulse on a single photodiode readout, and an accumulate-and-reset mode to test non-Gaussianity by attempting to surpass the Gaussian bound on an integrating photodiode readout. The former draws on existing sub-shot-noise demonstrations with photodiode readout. At the same time, the latter represents a realistic yet experimentally demanding next step, given the loss sensitivity and state preparation at modest mean photon numbers. Subject to standard linear-response modelling of the detector, an experimental violation of the detection-propagated Gaussian bound at the harvester would constitute an operational quantum advantage for precision charging—namely, a performance unattainable by any classical or Gaussian state under the same energy and temperature constraints. Together, these pathways make our precision-charging witnesses testable with state-of-the-art components and realistic losses, and they point to plausible near-term use cases, such as the precision charging of fragile nanoscopic loads (e.g., photodiode or quantum-dot elements), where reducing over-threshold events at a fixed mean delivered energy is valuable.
Beyond the single-mode analysis, we explicitly consider multimode batteries and entangling interferometers. For Gaussian resources, any passive linear network (beam splitters and phase shifters) that entangles the modes leaves the SNR $\Gamma$ invariant at fixed total energy and temperature (see Appendix G), whereas in the non-Gaussian regime passive mixing redistributes moments and can strictly increase $\Gamma$. This yields an interferometric optimization procedure and suggests a battery-side route to entanglement witnessing based solely on energy observables. We expect these multimode and entanglement-oriented extensions to broaden the operational scope of precision charging and to guide near-term experiments harnessing non-Gaussian resources on realistic, lossy platforms.
Acknowledgements
@@
The authors thank Antonis Delakouras, Manuel Gessner and Francesco Flora for fruitful discussions and inputs, and notably Antonio Acín for his relevant feedback and guidance.
Author contributions
B.P. developed the main research ideas, carried out the main calculations and simulations. F.C. proposed the main ideas and supervised the work. All authors reviewed and edited the manuscript.
Funding declaration
@@
This project has been funded by the European Union (EQC, 101149233), the Government of Spain (Severo Ochoa CEX2019-000910-S and FUNQIP), Fundació Cellex, Fundació Mir-Puig, Generalitat de Catalunya (CERCA program). This project is supported by the predoctoral program FI-STEP (2025 STEP 00008), which is backed by the Secretariat of Universities and Research of the Department of Research and Universities of the Generalitat of Catalonia, as well as the European Social Plus Fund. Views and opinions expressed are however, those of the author(s) only and do not necessarily reflect those of the European Union or European Research Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
Competing interests
The Authors declare no Competing Financial or Non- Financial Interests.
Appendix A Appendix A: Covariance matrix formalism for Gaussian states
@@
a squeezing transformation of parameter $z$. This representation allows interpreting every Gaussian state as the result of applying a single-mode squeezing transformation, followed by a linear-optical passive setup, to an initial thermal (and possibly displaced) state of $N$ modes. Such initial state would have covariance matrix $\sigma_{0}=\bigoplus_{j=1}^{N}\nu_{j}\mathds{1}_{2}$, where the $\nu_{i}$ are referred to as symplectic eigenvalues, and can be understood in this context as fluctuation parameters indicating the temperature $T_{i}$ that characterizes the initial thermal state of each of the modes through the relation:
[…]
The matrices $O_{1},O_{2}$ represent the energy-preserving passive optics transformations, which can be decomposed into beam splitters (semi-reflective mirrors parametrized by their transmissivity $\cos^{2}\theta$ that mix up pairs of modes) and phase shifters (dielectric plates that rotate the optical phase of an electromagnetic wave by some angle $\varphi$). A scheme for alternatingly applying these two operations to construct the most general linear interferometer is detailed in [TAN2019100030]. Matrix $Z$ represents local or single-mode squeezing, which, in contrast, is an active operation, requiring energy from an external laser field that pumps a nonlinear crystal through which the optical modes travel. The result of this transformation is a contraction of the variance of one of the canonical variables and the expansion of its conjugate one. Squeezing and beam splitting enable the generation of quantum correlations between different modes of a quantum state.
@@
$\sigma=\nu\begin{pmatrix}z\cos^{2}\varphi+\frac{1}{z}\sin^{2}\varphi&\frac{1-z^{2}}{2z}\sin(2\varphi)\\ \\ \frac{1-z^{2}}{2z}\sin(2\varphi)&z\sin^{2}\varphi+\frac{1}{z}\cos^{2}\varphi\end{pmatrix}$ (24)
The mean and variance of the photon number operator can be obtained through (see [Friis_2018])
@@
$\begin{split}(\Delta N(\rho))^{2}=\frac{1}{8}\nu^{2}\bigg(z^{2}+\frac{1}{z^{2}}\bigg)-\frac{1}{4}+\nu\bigg[\alpha_{1}^{2}(z\cos^{2}\varphi+\frac{1}{z}\sin^{2}\varphi)+\alpha_{2}^{2}(z\sin^{2}\varphi+\frac{1}{z}\cos^{2}\varphi)+2\alpha_{1}\alpha_{2}\bigg(\frac{1-z^{2}}{z}\sin\varphi\ \cos\varphi\bigg)\bigg]\end{split}$ (28)
@@
and through the quotient between Eqs.27 and 30 one can compute the ratio $\Gamma$ of any Gaussian state as a function of the squeezing and displacement transformations.
Appendix B Appendix B: Optimal signal-to-noise ratio $\Gamma$ for Gaussian states
@@
and finally the optimal Gaussian ratios $\Gamma$, which are shown in Fig.2 in the main text, are simply calculated by substituting these optimal solutions for $z$ and $\vec{\alpha}$ into Eqs.27 and 30.
To get a hint on how much of each energetic resource (squeezing and displacement) contributes to the optimal Gaussian state, we show the numerical values of the solutions as a function of $T,\epsilon$ in Fig.5. Whereas the optimal displacement appears to increase with increasing energy input to the system, regardless of temperature, the optimal squeezing shows a less uniform behaviour. Interestingly, for most of the $(T,\epsilon)$ values, highly squeezed states (i.e. $z_{opt}$ small) are advantageous, and so for the majority of the parameter space we may approximate
@@
Figure 5: Solution to 31 for the optimal squeezing and displacement through the Lagrange multipliers method.
@@
Appendix C Appendix C: Expectation values on non-Gaussian states